hallyramos12:

The area of a rectangle is 60 square inches. The length of the rectangle is 7 inches longer than the width. Which equation models this situation? w + 7w = 60 7w = 60 w + 7 = 60 w(w+7)=60

hallyramos12:

w( w + 7 ) = 60

hallyramos12:

I think

MrMudd183:

if you use that one, you will end up with \[w^2+7w=60\] this solution would leave you with more work to do. i think it's the first one

MrMudd183:

cause when you add w and 7w, you would get 8w then you divide both sides by 8 to get w by itself

hallyramos12:

oh

hallyramos12:

You have a point

MrMudd183:

after you dived both sides by 8, you should end up with w=7.5 if my math is correct

Mercury:

@hallyramos12 I agree with your original answer. if w = width, then "The length of the rectangle is 7 inches longer than the width." means that l = w + 7 length*width = area ---> w(w+7) = 60 which is in agreement with choice 4